OpenStudy (anonymous):

What is the value of the expression 2/4^-3 ? A. 1/32 B. 1/6 C. 128 D. 256

OpenStudy (anonymous):

Will Fan and give Medal.

OpenStudy (jango_in_dtown):

2/4^-3=2/2^-6=2^7=128

OpenStudy (jango_in_dtown):

C. is correct

OpenStudy (jango_in_dtown):

@xOmqKuro

OpenStudy (anonymous):

Oh, thanks!